OpenStudy (anonymous):

If (x + 2) is a factor of x3 − 6x2 + kx + 10 , k = please help !

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i have a suggestion!

OpenStudy (misty1212):

if \((x+2)\) is a factor, then if you replace \(x\) by \(-2\) you have to get zero

OpenStudy (misty1212):

put \(x=-2\) in to \(x^3+6x^2+kx+10=0\) and solve for \(k\)

OpenStudy (anonymous):

i thank that means put \[(-2)^3+6(-2)^2-2k+10=0\] right?

OpenStudy (misty1212):

that should do it you good from there?

OpenStudy (anonymous):

Yeah I should be good from here , but would you mind checking my answer when I'm done ?

OpenStudy (misty1212):

no not at all go ahead and post it

OpenStudy (anonymous):

So I just did it and the answer I got is -11

OpenStudy (misty1212):

oops

OpenStudy (misty1212):

\[-8+12-2k=10=0\] is a start

OpenStudy (anonymous):

Ok so I redid it and this time I got 2 as an answer

OpenStudy (misty1212):

yeah i made a mistake too

OpenStudy (misty1212):

first step is \[-8+24-2k+10=0\]

OpenStudy (misty1212):

then \[26-2k=0\\ 26=2k\\ 13=k\]

OpenStudy (anonymous):

Yeah I got 13 over here as well

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

thankyou so much for the help :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]